# What LLMs Miss in Recommendations: Bridging the Gap with Retrieval-Augmented Collaborative Signals

### v2

@@ Abstract.
User-item interactions contain rich collaborative signals that form the backbone of many successful recommender systems. While recent work has explored the use of large language models (LLMs) for recommendation, it remains unclear whether LLMs can effectively reason over this type of collaborative information. In this paper, we conduct a systematic comparison between LLMs and classical matrix factorization (MF) models to assess LLMs’ ability to leverage user-item interaction data. We further introduce a simple retrieval-augmented generation (RAG) method that enhances LLMs by grounding their predictions in structured interaction data. Our experiments reveal that current LLMs often fall short in capturing collaborative patterns inherent to MF models, but that our RAG-based approach substantially improves recommendation quality—highlighting a promising direction for future LLM-based recommenders.
@@ 1. Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) have demonstrated impressive capabilities in a wide range of tasks, including reasoning over textual input (Huang and Chang, 2022; Plaat et al., 2024), answering complex questions (Singhal et al., 2025; Jiang et al., 2021), generating fluent text (Gao et al., 2023), and encoding world knowledge (Singhal et al., 2023). As a result, LLMs have been increasingly adopted in diverse research domains such as data collection (Wei et al., 2024), summarization (Basyal and Sanghvi, 2023; Liu et al., 2023), translation (Zhu et al., 2023; Kocmi et al., 2024; Raunak et al., 2023), and visualization (Li et al., 2024; Wu et al., 2024a).
More recently, researchers have begun exploring the potential of LLMs in recommender systems, where the goal is to deliver personalized item suggestions that align with a user’s preferences—such as recommending movies, products, or articles (Bao et al., 2023; Dai et al., 2023; Sanner et al., 2023; Yue et al., 2023). This emerging line of work, often referred to as LLMs as recommenders (LLMRec), presents an exciting new research direction that leverages the capabilities of LLMs to tackle core challenges in recommender systems (Zhao et al., 2024; Bao et al., 2023; Zhang et al., 2023). Within this domain, LLMs have primarily been applied to two key subtasks: item understanding and user modeling. For example, in content-based filtering, LLMs can generate rich item representations from textual descriptions and infer user preferences from their interaction history (Lin et al., 2025; Wu et al., 2024b).
@@
One limitation of LLMRec methods is their insufficient modeling of collaborative information embedded in user-item co-occurrence patterns. To address this, Sun et al. (2024) proposed an approach that distills the world knowledge and reasoning capabilities of LLMs into a collaborative filtering recommender system. Their method adopts an in-context, chain-of-thought prompting strategy, focusing on the LLM-enhanced RS paradigm.
However, the integration of collaborative filtering within the LLMs-as-RSs paradigm—where recommendations are derived from patterns across many users—remains relatively underexplored. This gap is due in part to fundamental challenges: it is unclear how to adapt LLMs to effectively reason over collaborative signals, and the input length limitations of LLMs make it difficult to encode dense interaction histories needed to learn from similar users at scale.
In this paper, we take a step toward understanding this gap by analyzing the reasoning capabilities of LLMs over collaborative signals. Our first research question is: Can LLMs capture and reason over collaborative patterns? To assess the basic reasoning ability of LLMs in utilizing collaborative information, we compare their performance with matrix factorization, a classical and widely used method for modeling collaborative information in recommender systems. This comparison allows us to evaluate whether LLMs can reason over collaborative signals at a level comparable to—even if not superior to—a simple yet effective baseline like matrix factorization.
Our second research question investigates how to improve LLMs’ ability to reason over collaborative data. To this end and overcome to the challenges that mentioned above, we explore a retrieval-augmented generation (RAG) approach that supplies the LLM with relevant user-item interaction information at inference time, aiming to bridge the gap between LLMs and traditional collaborative filtering methods.
@@ 2. Preliminaries
In this study, we explore the reasoning ability of LLMs on collaborative filtering information. Our investigation is centered around a matrix of interactions between users and items, i.e., $M$. Let $\mathcal{U}=\{u_{1},\dots,u_{n}\}$ denote the set of all $n$ users and =$\mathcal{I}=\{v_{1},\dots,v_{m}\}$ denote the set of all $m$ items (movies). The matrix $M\in\mathbb{R}^{n\times m}$ represents the rating behavior of users over items, where each row corresponds to a user in $\mathcal{U}$ and each column to an item in $\mathcal{I}$. The entry $M_{ij}$ indicates the rating that user $u_{i}\in\mathcal{U}$ has given to item $v_{j}\in\mathcal{I}$, with values typically ranging from $1$ (disliked) to $5$ (highly liked), or 00 if the user has not rated the item.
To enable large language models (LLMs) to understand collaborative signals in the interaction matrix, it is necessary to provide them with structured information derived from the matrix. However, including the entire matrix in the input is infeasible, especially in large-scale datasets with many users and items. Moreover, presenting all such information in a prompt can overwhelm the LLM, making it difficult to interpret the collaborative structure and respond effectively to recommendation queries.
Recently, retrieval-augmented generation (RAG) has demonstrated impressive results in enhancing LLM performance across a variety of tasks (Fan et al., 2024; Chen et al., 2024; Zhang et al., 2024). RAG approaches integrate external retrieval mechanisms to supplement LLMs with relevant and targeted information, thereby improving the accuracy, relevance, and contextual understanding of generated responses.
In the following section, we present our method for applying retrieval-augmented generation (RAG) to enhance the effectiveness and efficiency of LLMs in leveraging collaborative filtering information.
@@ 3. Methodology
Given a target user $u_{t}$, our goal is to extract relevant signals from the interaction matrix $M$ to enable an LLM to predict and recommend movies that the user is likely to like.
To identify useful information, we first retrieve users most similar to the target user. These similar users are assumed to carry the most informative signals for inferring the preferences of $u_{t}$, as their past ratings can help guide the recommendation of unseen items.
To identify users most similar to the target user $u_{t}$, we construct a user–item interaction matrix by pivoting the masked dataset (test data are masked for all models) such that each row corresponds to a user and each column to an item, with missing ratings filled with zero. We then compute pairwise user similarities using cosine similarity:
@@
We define the set of top-$k$ similar users as:
@@ 3.1. Prompt Generation Strategies
We explore four distinct strategies for constructing our prompts. Each approach presents user rating data differently to investigate its impact on model performance and prompt efficiency.
@@ 3.1.1. Unfiltered Full Ratings Prompt (Baseline)
This baseline approach includes the full rating history of all top-$k$ similar users without applying any filtering or deduplication. Each user’s ratings are presented in their raw form. While the fraction parameter ($f$) still controls how much of the similar users’ data is included in the prompt, no preference is given to unseen items or highly rated content. This method offers a comprehensive view of all available preferences, potentially enriching the context provided to the LLM.
Notably, we include all sampled ratings from similar users, even for items that the target user has already rated. We refer to this approach as the baseline.
Example:
Target user A has rated the following movies: M1 (5), M2 (3), M3 (4)…
Top-$k$ similar users have rated:
User 101 rated: M1 (1), M7 (3), M2 (2)
User 24 rated: M2 (3), M4 (4)
Which 10 movies should user A watch next that they haven’t seen?
@@ 3.1.2. Sentiment-Based Prompt
This strategy organizes the ratings from similar users into three sentiment categories: Liked, Neutral, and Disliked. These categories are defined by rating thresholds: ratings $\geq 4$ are classified as Liked, a rating of $3$ as Neutral, and ratings $\leq 2$ as Disliked. As with other strategies, we exclude any movies that the target user has already rated to focus only on unseen content.
Example:
Target user A has rated the following movies: M1 (5), M2 (3), M3 (4)…
Top-$k$ similar users to user A have collectively rated the following unseen movies:
Liked: 234, 567, 890
Neutral: 111, 222
Disliked: 333, 444
Based on these patterns, which 10 movies should user A watch next?
@@ 3.1.3. Reasoning-Based Prompt
This strategy builds upon the baseline by introducing two key modifications. First, we remove any movies from the similar users’ ratings that have already been seen by the target user, ensuring that all recommended items are truly unseen. Second, we sample a fraction $f$ of the remaining ratings for inclusion in the prompt, controlling the prompt size while preserving informative signals.
Additionally, the prompt is explicitly framed as a reasoning task by appending a directive such as: “Reason based on the patterns above: which 10 movies should user A watch next that they haven’t seen?” This phrasing encourages the model to engage in logical inference rather than relying purely on pattern matching, potentially improving generalization and recommendation quality.
Example:
Target user A has rated the following movies: M1 (5), M2 (3), M3 (4)…
Top-$k$ similar users to user A have collectively rated the following unseen movies:
User 101 rated: M7 (3), M6 (2)
User 24 rated: M8 (3), M4 (5)
Reason based on the patterns above: which 10 movies should user A watch next that they haven’t seen?
@@ 3.1.4. Full Reasoning
In the Reasoning approach, the LLM only sees data from the top-$k$ similar users, and therefore lacks access to global signals—such as the overall popularity of movies across the user base. To address this limitation, we propose an enhanced variant that augments the prompt with additional popularity-related statistics.
For each movie $v_{j}\in\mathcal{I}$, we compute two global metrics based on the interaction matrix $M\in\mathbb{R}^{n\times m}$:
  • Rating Count: The number of users who rated movie $v_{j}$:
@@
  • Average Rating: The average rating received by movie $v_{j}$:
@@
These statistics are appended to each corresponding movie entry in the ”Reasoning” prompt. This allows the LLM to consider not only the preferences of similar users but also the broader appeal and quality of each recommended item.
Example:
Target user A has rated the following movies: M1 (5), M2 (3), M3 (4)…
Top-$k$ similar users to user A have collectively rated the following unseen movies:
User 101 rated: M7 (3), M6 (2)
User 24 rated: M8 (3), M4 (5)
Movie popularity stats:
M4 — Count: 231, AvgRating: 4.2
M6 — Count: 121, AvgRating: 3.6
M7 — Count: 298, AvgRating: 3.8
M8 — Count: 83, AvgRating: 3.2
Reason based on the patterns above and the popularity statistics: which 10 movies should user A watch next that they haven’t seen?
@@ 4. Evaluation
Figure 2. NDCG score for different prompt generation strategies and MF as a function of number of similar users ($k$) and different percentage of their information in the prompt ($f$). Number of users is $350$ in both cold and hot groups. Top is the hot users and bottom is cold users.
@@ 4.1. Dataset
We evaluate the effectiveness of using LLMs for personalized movie recommendation using the widely studied MovieLens 100K dataset. This dataset comprises $100,000$ ratings from $943$ users on $1,682$ movies. As a preprocessing step, we remap both user and movie IDs to be contiguous integers starting from zero to ensure consistency across model input formats. We sort each user’s rating history chronologically and randomly mask 20% of their ratings to simulate unobserved preferences. These masked ratings are withheld and later used to evaluate recommendation accuracy, while the remaining 80% of ratings are retained to simulate user history and compute user-user similarities.
To evaluate the models fairly across different user profiles, we split the users into two groups based on activity level. Users with a number of ratings above the dataset’s median are categorized as ”hot users,” and the rest as ”cold users.” We randomly sample $350$ users from each group.
@@ 4.2. Experimental and parameter settings
For each user, we construct a prompt that includes (1) the known ratings from their retained history, and (2) the ratings from the top-k most similar users. These similar users are selected using cosine similarity computed over the masked user-item matrix. To simulate different levels of auxiliary knowledge, we vary the fraction of each similar user’s ratings shown in the prompt from 25% to 100%. We use $f$ to indicate the fraction value. We also explicitly inform the LLM about the range of movie IDs in the dataset to provide context.
The prompt concludes with a question asking the model to suggest $10$ movies that the target user has not seen. The LLM is then queried using OpenAI’s ChatCompletion API, and its response is parsed to extract movie IDs. These predictions are evaluated using NDCG (Normalized Discounted Cumulative Gain) and Hit@10 metrics. NDCG evaluates how well a ranked list puts the most relevant items near the top and Hit@10 metric measures the proportion of ground-truth masked movies that appear in the model’s top-10 recommendations. Formally, for each user, Hit@10 is calculated as the cardinality of the intersection between the predicted and held-out movies divided by $10$. We compute and average this value across all users in both the hot and cold user groups.
Figure 3. Hit@10 score for different prompt generation strategies and MF as a function of number of similar users ($k$) and different percentage of their information in the prompt ($f$). Number of users is $350$ in both cold and hot groups. Top is the hot users and bottom is cold users.
@@
To establish a strong baseline, we also implement a matrix factorization model in PyTorch and train it using stochastic gradient descent over a range of hyperparameters, including different numbers of latent factors ({10, 20, 50, 100}) and batch sizes ({8, 16, 32, 64,128, and 256}). The model is trained for up to $3,000$ epochs or until convergence, with performance evaluated on a held-out test set. We report the best-performing configuration after convergence based on the highest NDCG score for the testing cases. The model is trained on the same 80% unmasked ratings used in the LLM setup.
4.3. Results and Discussion
Figure 2 presents the average NDCG scores obtained across different prompt generation strategies as a function of the number of similar users $k$, for multiple fractions of their rating histories ($f$). We observe that reasoning-based prompts (such as reasoning and full reasoning) consistently improve in performance as $k$ increases. This suggests that LLMs benefit from richer collaborative information when it is presented in a structured and interpretable format. In contrast, simpler strategies such as sentiment-based or original tend to decrease in NDCG as $k$ increases. This likely occurs because these prompts include more raw data without structure or abstraction, making it harder for the LLM to extract relevant signals—highlighting the importance of structured prompt engineering. These results support integrating techniques like retrieval-augmented generation (RAG) or knowledge graphs to organize collaborative user–item interactions more effectively.
As we increase the fraction of similar user data included in the prompt, we observe a general increase in overall NDCG scores across reasoning strategies. This aligns with the expectation that providing more evidence to the LLM leads to stronger predictions. However, an interesting trend appears when comparing cold vs. hot users: LLMs tend to perform better on cold users than hot users, while Matrix Factorization (MF) exhibits the opposite behavior—achieving higher scores for hot users and lower for cold ones. This is expected, as MF relies on sufficient user-item interactions to learn embeddings, making it more effective for users with rich histories.
Figure 4 shows the cumulative distribution function (CDF) of prompt processing times for cold and hot users using the reasoning_ranked prompt generation strategy, with the number of similar users fixed at $k=10$, and varying the fraction of each similar user’s rating history included in the prompt. By processing time, we refer to the total latency involved in communicating with the GPT server, including the time to transmit the prompt, the model’s internal computation, and the time required to receive the response back to the local machine.
For cold users, the CDF curves for different fraction values (e.g., $0.25,0.5,0.75,1.0$) are tightly clustered and exhibit nearly identical processing times. This is expected because cold users, by definition, have rated very few movies. Consequently, even when using the full fraction of ratings from their similar users, the total number of ratings included in the prompt remains small. Thus, the prompt length—and hence the request latency—does not vary significantly with the fraction parameter.
In contrast, for hot users, we observe that higher fraction values lead to noticeably longer processing times. Since hot users have many ratings, each similar user is likely to contribute more movie–rating pairs as the fraction increases. This results in significantly longer prompts, which take longer to serialize, transmit to the model, and process on the server. The effect is visible in the rightward shift of the CDF curves for higher fractions.
This analysis highlights that prompt length has a strong effect on processing latency, especially for users with rich interaction histories. It also supports the importance of designing token-efficient prompt generation strategies when scaling LLM-based recommendation systems, particularly for hot users with high $k$ and large data fractions.

### v3

# Do LLMs Understand Collaborative Signals? Diagnosis and Repair

@@ Abstract.
Collaborative information from user-item interactions is a fundamental source of signal in successful recommender systems. Recently, researchers have attempted to incorporate this knowledge into large language model-based recommender approaches (LLMRec) to enhance their performance. However, there has been little fundamental analysis of whether LLMs can effectively reason over collaborative information. In this paper, we analyze the ability of LLMs to reason about collaborative information in recommendation tasks, comparing their performance to traditional matrix factorization (MF) models. We propose a simple and effective method to improve LLMs’ reasoning capabilities using retrieval-augmented generation (RAG) over the user-item interaction matrix with four different prompting strategies. Our results show that the LLM outperforms the MF model whenever we provide relevant information in a clear and easy-to-follow format, and prompt the LLM to reason based on it. We observe that with this strategy, in almost all cases, the more information we provide, the better the LLM performs.
@@ 1. Introduction
Large Language Models (LLMs) (achiam2023gpt; touvron2023llama) have demonstrated impressive capabilities in a wide range of tasks, including reasoning over textual input (huang2022towards; plaat2024reasoning), answering complex questions (singhal2025toward; jiang2021can), generating fluent text (gao2023enabling), and encoding world knowledge (singhal2023large). As a result, LLMs have been increasingly adopted in diverse research domains such as data collection (wei2024llmrec), summarization (basyal2023text; liu2023learning), translation (zhu2023multilingual; kocmi2024findings; raunak2023leveraging), and visualization (li2024visualization; wu2024automated).
More recently, researchers have begun exploring the potential of LLMs in recommender systems, where the goal is to deliver personalized item suggestions that align with a user’s preferences—such as recommending movies, products, or articles (bao2023tallrec; dai2023uncovering; sanner2023large; yue2023llamarec). This emerging line of work, often referred to as LLMs as recommenders (LLMRec), presents an exciting new research direction that leverages the capabilities of LLMs to tackle core challenges in recommender systems (zhao2024recommender; bao2023tallrec; zhang2023chatgpt). Within this domain, LLMs have primarily been applied to two key subtasks: item understanding and user modeling. For example, in content-based filtering, LLMs can generate rich item representations from textual descriptions and infer user preferences from their interaction history (lin2025can; wu2024survey; liao2024llara).
@@
One limitation of LLMRec methods is their insufficient modeling of collaborative information embedded in user-item co-occurrence patterns. To address this, sun2024large proposed an approach that distills the world knowledge and reasoning capabilities of LLMs into a collaborative filtering recommender system. Their method adopts an in-context, chain-of-thought prompting strategy, focusing on the LLM-enhanced RS paradigm.
However, the ability of LLMs to reason over collaborative filtering information remains limited. Recent work (e.g., (kim2024large)) attempts to bridge this gap by aligning user–item embeddings with the LLM’s representation space, but such methods still fall short—LLMs cannot directly access or reason over the full set of similar users’ interactions due to input and context constraints.
In this paper, we take a step toward addressing this gap by analyzing the reasoning capabilities of LLMs over collaborative signals. Our first research question is: Can LLMs capture and reason over collaborative patterns when provided with necessary information, such as similar users or items? To examine this, we compare LLMs with matrix factorization—a classical and widely used method for modeling collaborative information in recommender systems. This comparison enables us to assess whether LLMs can reason over collaborative signals at a level comparable to this simple yet strong baseline.
Our second research question examines how to enhance LLMs’ ability to reason over collaborative data. To address the challenges discussed above, we explore a retrieval-augmented generation (RAG) approach that provides the LLM with relevant user–item interaction information at inference time, aiming to bridge the gap between LLMs and traditional collaborative filtering methods. With this strategy, we reduce noise in the LLM’s input and help it reason with greater precision.
We examine four different prompting strategies and show that providing relevant information in a clear format and prompting the LLM to reason over it significantly improves performance, especially for cold-user data. We also show that, given relevant information, the LLM can effectively continue making recommendations as more context becomes available—for example, from additional similar users or more items associated with them.
@@ 2. Preliminaries
In this study, we explore the reasoning ability of LLMs on collaborative filtering information. Our investigation is centered around a matrix of interactions between users and items, i.e., $M$. Let $\mathcal{U}=\{u_{1},\dots,u_{n}\}$ denote the set of all $n$ users and $\mathcal{I}=\{v_{1},\dots,v_{m}\}$ denote the set of all $m$ items (movies). The matrix $M\in\mathbb{R}^{n\times m}$ represents the rating behavior of users over items, where each row corresponds to a user in $\mathcal{U}$ and each column to an item in $\mathcal{I}$ as shown in the left side matrix in figure 1. The entry $M_{ij}$ indicates the rating that user $u_{i}\in\mathcal{U}$ has given to item $v_{j}\in\mathcal{I}$, with values typically ranging from $1$ (disliked) to $5$ (highly liked), or 0 if the user has not rated the item.
To enable LLMs to understand collaborative signals in the interaction matrix, it is necessary to provide them with structured information derived from the matrix. However, including the entire matrix in the input is infeasible, especially in large-scale datasets with many users and items. Moreover, presenting all such information in a prompt can overwhelm the LLM, making it difficult to interpret the collaborative structure and respond effectively to recommendation queries.
Recently, retrieval-augmented generation (RAG) has demonstrated impressive results in enhancing LLM performance across a variety of tasks (fan2024survey; chen2024benchmarking; zhang2024r). RAG approaches integrate external retrieval mechanisms to supplement LLMs with relevant and targeted information, thereby improving the accuracy, relevance, and contextual understanding of generated responses.
In the following section, we present our method for applying RAG to enhance the effectiveness and efficiency of LLMs in leveraging collaborative filtering information.
@@ 3. Methodology
Given a target user $u_{t}$, our goal is to extract relevant signals from the interaction matrix $M$ to enable an LLM to predict and recommend movies that the user is likely to like. To identify useful information, we first retrieve users most similar to the target user. These similar users are assumed to carry the most informative signals for inferring the preferences of $u_{t}$, as their past ratings can help guide the recommendation of unseen items.
To identify users most similar to the target user $u_{t}$, we compute pairwise user similarities using cosine similarity over their rating vectors: :
@@
where $M_{u_{t}}$ and $M_{u}$ denote the rating vectors of users $u_{t}$ and $u$, respectively. Each vector is constructed from the corresponding row in the interaction matrix $M$, with unobserved ratings treated as zeros. We define the set of top-$k$ similar users as:
@@ 3.1. Prompt Generation Strategies
We explore four distinct strategies for constructing our prompts. Each approach presents user rating data differently to investigate its impact on model performance and prompt efficiency. We explain this approach and provide one example of each strategy in figure 1.
@@ 3.1.1. Unfiltered Full Ratings Prompt (Baseline)
This baseline incorporates the complete rating histories of the top-$k$ most similar users without applying any filtering or prioritization. Each user’s ratings are included in their original, unprocessed form. Although the fraction parameter ($f$) still determines the proportion of similar users’ data included in the prompt, the selection does not favor unseen items or those with higher ratings. By presenting all available preference data, this approach aims to provide a broad and potentially richer context to the LLM.
Importantly, this approach includes all sampled ratings from similar users, even for items that the target user has already rated. This design choice allows the LLM to observe both the target user’s and similar users’ opinions on the same items, potentially helping it better infer the target user’s preferences. At the same time, we explicitly instruct the LLM not to select movies that the target user has already rated. We refer to this strategy as our baseline.
@@ 3.1.2. Sentiment-Based Prompt
In this strategy, ratings from similar users are grouped into three sentiment categories: Liked, Neutral, and Disliked. These categories are determined using rating thresholds: ratings $\geq 4$ are considered Liked, a rating of $3$ is treated as Neutral, and ratings $\leq 2$ fall under Disliked. Additionally, we exclude any movies that the target user has already rated to ensure the model focuses solely on unseen content.
Figure 2. NDCG score for different prompt generation strategies and MF as a function of $k$ and $f$.
@@ 3.1.3. Reasoning-Based Prompt
The interaction matrix in collaborative filtering captures valuable relational signals that can be interpreted as chains of reasoning, leading to meaningful recommendations. For example, if user A and user B both rated a particular item as 2, we can infer that their preferences are similar. If user B then gives a rating of 5 to another item, it is reasonable to predict that user A may also enjoy that item.
While this example involves just two users and two items, the underlying idea extends to more complex scenarios involving many users and items. This kind of transitive inference demonstrates how collaborative signals can be composed to uncover deeper and more nuanced patterns of user preferences.
Inspired by the success of chain-of-thought prompting for LLMs (wei2022chain), we design this strategy to encourage structured reasoning over such collaborative chains. Specifically, we build on the baseline by introducing two key modifications. First, we exclude any movies from similar users’ ratings that have already been rated by the target user, ensuring that all recommended items are truly unseen. Second, we explicitly frame the prompt as a reasoning task by appending a directive such as: “Reason based on the patterns above: which 10 movies should user A watch next that they haven’t seen?” This formulation encourages the LLM to perform inference based on user-item interaction patterns, rather than relying purely on shallow matching.
By structuring the task in this way, we implicitly apply a “step-by-step” reasoning strategy, which has been shown to improve the performance of LLMs across a range of tasks (wei2022chain).
@@ 3.1.4. Full Reasoning
In the Reasoning approach, the LLM is exposed only to the preferences of the top-$k$ most similar users and therefore lacks access to global information—such as how popular a movie is across the entire user base. To address this limitation, we introduce an enhanced variant that augments the prompt with additional popularity-based statistics.
For each movie $v_{j}\in\mathcal{I}$, we compute two global metrics derived from the interaction matrix $M\in\mathbb{R}^{n\times m}$:
  • Rating Count: The number of users who have rated movie $v_{j}$:
@@
  • Average Rating: The mean rating received by movie $v_{j}$:
@@
These statistics are appended to each corresponding movie entry in the Reasoning prompt. By doing so, the LLM is equipped to weigh both local signals (from similar users) and global signals (reflecting overall popularity and quality), potentially improving the relevance and diversity of its recommendations.
@@ 4.1. Dataset
We evaluate our methods and baselines using the widely studied MovieLens 100K dataset (harper2015movielens), which contains 100,000 ratings provided by 943 users on 1,682 movies. As part of preprocessing, we remap both user and movie IDs to contiguous integers starting from zero to ensure compatibility with our model input formats.
To simulate realistic recommendation scenarios, we sort each user’s rating history chronologically and randomly mask 20% of their ratings to represent unobserved preferences. These masked ratings are withheld during training and later used for evaluation. The remaining 80% of the ratings are treated as known history and are used to compute user-user similarities and construct prompts for the language model.
Figure 3. Hit@10 score for different prompt generation strategies and MF as a function of $k$ and $f$.
To ensure fair evaluation across varying user profiles, we partition users into two groups based on their activity level. Specifically, users with a number of ratings above the dataset’s median are labeled as hot users, while those below or equal to the median are considered cold users. We randomly sample 350 users from each group to construct balanced evaluation sets.
@@ 4.2. Experimental and parameter settings
To simulate varying levels of auxiliary knowledge, we control the fraction $f$ (25% to 100%) of each similar user’s ratings included in the prompt, allowing us to examine how contextual information impacts recommendation quality. We also inform the LLM of the valid movie ID range to help it interpret the data consistently. Each prompt ends by asking the model to recommend 10 unseen movies for the target user. We query the GPT-4.1-mini model via OpenAI’s ChatCompletion API and extract recommended movie IDs from its response. These are evaluated using Normalized Discounted Cumulative Gain (NDCG) (jarvelin2002cumulated), which rewards correctly ranked relevant items, and Hit@10, which measures how often ground-truth masked movies appear in the top 10 results.
4.3. Results and Discussion
Figures 2 and 3 show average NDCG and Hit@10 scores across prompt generation strategies, varying the number of similar users $k$ and fraction $f$ of their ratings included. Reasoning-based prompts consistently improve with larger $k$, indicating LLMs benefit from structured, interpretable collaborative signals. In contrast, simpler strategies (e.g., sentiment-based or original) decline as $k$ increases, likely due to unstructured raw data overwhelming the model. These findings underscore the need for structured prompt design, potentially leveraging RAG or knowledge graphs.
As the fraction of similar user data in the prompt increases, NDCG scores improve for reasoning strategies, confirming that more context helps LLM predictions. Notably, LLMs perform better on cold users, while Matrix Factorization (MF) favors hot users due to its reliance on rich interaction histories for learning effective embeddings.
@@ 4.3. Results and Discussion
Figure 4 shows the CDF of prompt processing times and prompt sizes (number of tokens) for cold and hot users using the Full reasoning strategy, with $k=10$ and varying fractions of each similar user’s ratings. Processing time includes prompt transmission, GPT computation, and response retrieval. For cold users, CDF curves across fractions (0.251.00.25-1.0) are nearly identical, as they have few ratings and adding more from similar users does not significantly increase prompt length or latency. While increasing the fraction does slightly increase token count for cold users, their prompts remain shorter than those of hot users. In contrast, hot users show clear increases in processing time with higher fractions-larger histories lead to longer prompts and higher latency, evident from the rightward shift in the CDF curves. These findings highlight that prompt length strongly impacts latency for hot users and emphasize the need for token-efficient prompt strategies in scalable LLM-based recommendation systems.